St. Teresa's Medical Center (STMC) offers a number
Answer #1
Costs Patients Days Cost Beaviour Remark
4200 4500
Salaries Nurse 55000 55000 Fixed Cost Since the Total Cost Remained Same in two different units of operation.
Aides 32000 32000 Fixed Cost Since the Total Cost Remained Same in two different units of operation.
Pharmacy 235700 251300 Mixed Cost A portion of Cost is fixed and and the Remaining portion is Variable
Laboratory 120300 127200 Mixed Cost A portion of Cost is fixed and and the Remaining portion is Variable
Depreciation 25000 25000 Fixed Cost Since the Total Cost Remained Same in two different units of operation.
Laundry 20160 20160 Fixed Cost Since the Total Cost Remained Same in two different units of operation.
Administration 27000 27000 Fixed Cost Since the Total Cost Remained Same in two different units of operation.
Lease (equipment) 36000 36000 Fixed Cost Since the Total Cost Remained Same in two different units of operation.
Per unit Calculation
Pharmacy
For 4200 Paitents Days 56.12
For 4500 Paitents Days 55.84
Per unit Calculation
Laboratory
For 4200 Paitents Days 28.64
For 4500 Paitents Days 28.27
Both the Costs are Mixed in Nature, A portion of Cost is fixed and and the Remaining portion is Variable
Units Pharmacy Laboratory
For 4500 Paitents Days 4500 251300 127200
For 4200 Paitents Days 4200 235700 120300
Difference 300 15600 6900
Variable Cost Per Unit = 15600/300 = 6900/300
52 23
Total Variable Cost for 4500 units 234000 103500
Fixed Cost (Total Cost -Variable Cost) 17300 23700
Cost for 4300 patients days
Costs Amounts
Salaries Nurse 55000
Aides 32000
Depreciation 25000
Laundry 20160
Administration 27000
Lease (equipment) 36000
Pharmacy -Fixed Portion 17300
Laboratory -Fixed Portion 23700
Total Fixed Cost 236160
Variable Cost
Pharmacy -Variable Portion (4300*52) 223600
Laboratory -Variable Portion (4300*23) 98900
Total Variable Cost 322500
Total Cost for 4300 patients Days 558660
Amount cargable per patients days 129.92
Cost for 4800 patients days
Costs Amounts
Salaries Nurse 55000
Aides 32000
Depreciation 25000
Laundry 20160
Administration 27000
Lease (equipment) 36000
Pharmacy -Fixed Portion 17300
Laboratory -Fixed Portion 23700
Total Fixed Cost 236160
Variable Cost
Pharmacy -Variable Portion (4800*52) 249600
Laboratory -Variable Portion (4800*23) 110400
Total Variable Cost 360000
Total Cost for 4800 patients Days 596160
Amount cargable per patients days 124.20